# Adaptive Accountability in Networked MAS:

### v1

@@ 7.3 Quantitative Results
Across all 360 configurations, AAF (i) reduces the compromise ratio from $C_{T}/T=0.48\pm 0.11$ (PPO-only) to $0.07\pm 0.02$, (ii) achieves a median detection delay of 9 steps (95th-percentile 17), in line with Corollary 2, and (iii) lifts mean social welfare by 12121818% over PPO-only while reducing the Gini index by up to 3333%. Bandwidth overhead never exceeds 5.4% of the application traffic budget.
Table 1 summarises the N=10,50N=\!10,50 slices. When $N=100$ (not shown for space) the trends persist but absolute rewards fall $\approx$10% because congestion dominates. A two-way ANOVA confirms significant effects of partialObs and penaltyFactor (interaction $p<.01$). High penalties without observability increase inequality, but enabling partial observability or a redistribution rule ($\alpha=0.25$) offsets this effect.
Table 1: Complete results for all parameter combinations (55 seeds per configuration).
Agents Penalty Partial Obs Dist. Policy Avg. Reward Final Gini
10 0.05 off alpha 81.20 0.117
10 0.05 off none 274.26 0.005
10 0.05 on alpha 135.11 0.001
10 0.05 on none 267.65 0.006
10 0.20 off alpha 110.09 0.136
10 0.20 off none 111.29 0.035
10 0.20 on alpha 114.25 0.103
10 0.20 on none 112.50 0.033
50 0.05 off alpha 15.39 0.032
50 0.05 off none 24.53 0.139
50 0.05 on alpha 178.01 0.020
50 0.05 on none 164.80 0.029
50 0.20 off alpha 130.29 0.019
50 0.20 off none 134.55 0.162
50 0.20 on alpha 239.59 0.011
50 0.20 on none 4.28 0.037
Figure 2 plots the final Gini by penalty factor and observability; error bars span seed variability. In almost all conditions the accountability layer nudges the Gini below 0.05—a level often cited as “near-egalitarian” in energy-grid studies [Rahwan2019MachineBehavior]. Outliers correspond to transient hoarding bursts that were detected but not fully countered before run end; these motivate future work on dynamic penalty scaling.
Figure 2: Final Gini by penalty factor ($x$-axis) and observability (colour). Error bars show ±1 s.d. across seeds.
@@ 7.4 Qualitative Case Study
In the $N=50$, penalty=0.05, partialObs=off, $\alpha=1$ scenario, a subset of four agents formed a “cartel” that requested $0.8R_{\max}$ each step. The governor flagged the resulting spike in Gini at $t=137$; reward shaping triggered at $t=146$, lowering the cartel’s average per-step gain from $10.4$ to $3.1$. Communication throttling then muted the highest-MI link; the cartel disintegrated by $t\approx 190$. Subsequent causal-graph visualization in the Grafana dashboard showed responsibility flowing back to previously sidelined agents, and the final compromise ratio settled at 0.08—well under the $\eta^{\star}=0.11$ analytical ceiling for that parameter set.
@@ 7.5 Reproducibility Checklist
  • Code & data: Public GitHub repository: https://github.com/alqithami/aaf-sim. Tag v1.0 reproduces all figures and tables.
  • Environment: flake.nix specifies exact library versions; Docker images published to GHCR under the same tag.
  • Randomness: All PRNG seeds recorded in ledger metadata and exported as seeds.csv.
  • Hardware: CPU/GPU specs in §6; no vendor-specific features required.
Empirical evidence aligns with the analytical bounds: AAF curbs harmful norms within nine steps median, keeps the compromise ratio below $\eta^{\star}$, and delivers double-digit gains in both welfare and equity—all while consuming $5\%$ additional bandwidth and CPU.
@@ 8.1 Why the Results Matter
Across every Monte-Carlo seed (§7) the compromise ratio fell below the target $\eta^{\star}$ within nine control steps median—roughly two seconds in our 5Hz traffic-signal benchmark. Classical “audit-once” schemes could not have reacted in time; by the time an offline audit completed, the resource pool would already have been captured. The finding supports the behavioral-science claim that timely sanctions deter norm-violating coalitions more effectively than harsher but delayed punishment [ostrom1990governing].
@@
The bounded-compromise theorem (Thm. 5) predicted that a local, Shapley-weighted penalty would globally stabilize the system once $\lambda H>g_{\max}$. Empirically, we observed $\eta_{\text{emp}}$ within one standard error of the analytic ceiling, even when 5% of agents behaved adversarially. This corroborates the long-standing conjecture that potential-based shaping [ng1999policy] scales from single- to multi-agent settings when augmented with causal attribution.
@@ 8.2 Limitations and Threats to Validity
Our grid spanned $\{10,50,100\}$ agents and two penalty magnitudes. A 5000-agent swarm, or penalty factors close to unity, might stress the message bus or saturate the intervention budget. Future work should stress-test the framework under orders-of-magnitude larger populations and adversaries who actively spoof causal edges.
@@ 8.3 Ethical, Legal, and Social Implications
Reward penalties can skew agents toward risk-averse behavior, potentially lowering innovation in financial-market MAS [ederer_manso_2013_p4p_innovation]. A human-in-the-loop override—triggered by the yellow-flag dashboard—helps balance safety with exploration.
@@
By design, Shapley-style attribution apportions blame according to marginal causal contribution. If agents differ in observability or computing power, this may correlate with protected attributes, leading to distributive inequity [mittelstadt2019principles]. Auditors should therefore scrutinize $\rho_{i}(e)$ distributions for disparate impact.
@@ 8.4 Relevance to Emerging AI Governance
The EU AI Act (2024) and the NIST AI Risk-Management Framework (2023) both highlight collective emergent risks —but offer few technical recipes for tracing responsibility in networked AI systems. AAF operationalizes these policy aspirations by (i) anchoring an immutable event ledger; (ii) providing real-time statistical tests with formal $\alpha$-level guarantees [howard2021time]; and (iii) linking those tests to proportionate, auditable interventions. Hence the framework can serve as a reference architecture for regulators drafting sector-specific codes of practice.
@@ 8.5 Future Directions
While the Grafana DAG explorer provides low-level proofs, policy operators may prefer natural-language rationales. Coupling the ledger to an LLM-based explainer guarded by the event digests to prevent hallucination could bridge the accountability-interpretability gap [kroll2017accountable].
@@ A.3 Proof of Theorem 3
Under $H_{0}$ the $m$-lag Granger residual–variance ratio has exact $F$ distribution F(m,nm)/mF(m,n{-}m)/m with n=m+1n=m+\!1 degrees-of-freedom [lutkepohl2005new, howard2020time]. Let $U_{t}\!=\!\mathbf{1}\{F_{t}>h_{t}\}$. Then, by the Chernoff bound for $\chi^{2}$ tails, $\mathbb{P}[U_{t}=1]\leq e^{-h_{t}^{2}/2}=e^{-h_{0}^{2}/2}\,t^{-1}.$ Taking the union bound over $t\geq 1$ yields
@@ A.4 Proof of Theorem 4
Define $f(h)=\mathbb{P}_{H_{0}}(S\geq h)-\alpha$. The process $\{h_{t}\}$ satisfies the Robbins-Monro recursion $h_{t+1}=h_{t}-\eta_{t}f(h_{t})+\eta_{t}\epsilon_{t+1}$ with martingale difference $\epsilon_{t+1}=Z_{t+1}-\mathbb{P}_{H_{0}}(S\geq h_{t})$. Because $S_{t}$ has bounded increments, $\epsilon_{t+1}$ is square integrable with $\mathbb{E}\,\epsilon_{t+1}=0$ and $\sup_{t}\mathbb{E}\,\epsilon_{t+1}^{2}\!<\infty$. Assumption $\sum_{t}\eta_{t}^{2}<\infty$ ensures the Kushner-Clark condition, so $h_{t}\!\to\!h^{\star}$ a.s. where $f(h^{\star})=0$ [kushner2003stochastic]. Ergodicity of the CUSUM statistic under $H_{0}$ implies $\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\{S_{t}\geq h^{\star}\}\to\mathbb{P}_{H_{0}}(S\geq h^{\star})=\alpha$, and dominated convergence gives the required limit for $h_{t}$. $\Box$

### v2

@@ 7.3 Quantitative Results
Grid-level effects (324 regimes).
Across the full factorial sweep (2 tasks $\times$ 3 scales $\times$ 3 penalties $\times$ 3 $\alpha$ values $\times$ 2 observability modes $\times$ 3 Byzantine rates), AAF-full reduces the executed compromise ratio relative to PPO-only in 96% of regimes, and this reduction is significant at $p<0.05$ in 86.4% of regimes (paired, seed-matched tests; 10 pairs per regime; Holm–Bonferroni corrected). The median relative reduction is 11.9%. Social welfare is preserved (median change 0.4%) and improves in 93% of regimes (significant in 78.1%).
Canonical baseline comparison (resource sharing, no Byzantines).
Table 1 reports a head-to-head comparison on resource_sharing under a canonical setting ($T=2000$, $\texttt{penalty\_factor}=0.2$, $\alpha=1.0$, full observability, $\rho=0$). AAF-full consistently reduces executed compromises relative to PPO-only across scales, with small welfare changes. Static-guard provides an oracle lower bound on executed compromise (zero by construction) but does not produce attribution signals.
Table 1: Baseline comparison on resource_sharing in the canonical setting ($T=2000$, $\texttt{penalty\_factor}=0.2$, $\alpha=1.0$, full observability, $\rho=0$). We report mean $\pm$ 95% CI over 10 seeds with 3 independent replications per seed ($n=30$ runs). Lower is better for Compromise and Gini; higher is better for Welfare.
NN Baseline Compromise ratio $\downarrow$ Social welfare $\uparrow$ Gini (Alloc.) $\downarrow$ Gini (Reward) $\downarrow$
10 AAF (Full) 0.411 $\pm$ 0.007 12.918 $\pm$ 0.001 0.203 $\pm$ 0.003 0.154 $\pm$ 0.003
PPO Only 0.520 $\pm$ 0.009 12.896 $\pm$ 0.002 0.206 $\pm$ 0.004 0.156 $\pm$ 0.003
Constrained PPO 0.519 $\pm$ 0.009 12.896 $\pm$ 0.002 0.206 $\pm$ 0.004 0.156 $\pm$ 0.003
Fair PPO 0.520 $\pm$ 0.009 12.896 $\pm$ 0.002 0.206 $\pm$ 0.004 0.156 $\pm$ 0.003
Static Guard 0.000 $\pm$ 0.000 13.000 $\pm$ 0.000 0.161 $\pm$ 0.007 0.124 $\pm$ 0.005
50 AAF (Full) 0.626 $\pm$ 0.005 2.475 $\pm$ 0.001 0.156 $\pm$ 0.002 0.109 $\pm$ 0.001
PPO Only 0.754 $\pm$ 0.007 2.449 $\pm$ 0.001 0.141 $\pm$ 0.002 0.103 $\pm$ 0.002
Constrained PPO 0.752 $\pm$ 0.007 2.450 $\pm$ 0.001 0.142 $\pm$ 0.003 0.104 $\pm$ 0.002
Fair PPO 0.754 $\pm$ 0.007 2.449 $\pm$ 0.001 0.141 $\pm$ 0.003 0.103 $\pm$ 0.002
Static Guard 0.000 $\pm$ 0.000 2.600 $\pm$ 0.000 0.103 $\pm$ 0.003 0.079 $\pm$ 0.003
100 AAF (Full) 0.634 $\pm$ 0.005 1.173 $\pm$ 0.001 0.167 $\pm$ 0.002 0.109 $\pm$ 0.001
PPO Only 0.706 $\pm$ 0.007 1.159 $\pm$ 0.001 0.159 $\pm$ 0.002 0.108 $\pm$ 0.001
Constrained PPO 0.700 $\pm$ 0.007 1.160 $\pm$ 0.001 0.161 $\pm$ 0.003 0.109 $\pm$ 0.002
Fair PPO 0.700 $\pm$ 0.007 1.160 $\pm$ 0.001 0.160 $\pm$ 0.002 0.108 $\pm$ 0.001
Static Guard 0.000 $\pm$ 0.000 1.300 $\pm$ 0.000 0.093 $\pm$ 0.003 0.072 $\pm$ 0.002
Ablation (what matters).
Table 2 isolates each AAF component. In this benchmark, patching accounts for the bulk of compromise reduction; detector-only and shaping-only variants largely match PPO-only on executed compromise, while patch-only recovers most of the safety gain.
Table 2: AAF ablation on resource_sharing under the same canonical setting as Table 1. Values are mean $\pm$ 95% CI over 10 seeds with 3 independent replications per seed ($n=30$ runs).
NN Variant Compromise ratio $\downarrow$ Social welfare $\uparrow$ Mean Gini (alloc.) $\downarrow$
10 AAF (Full) 0.411 $\pm$ 0.007 12.918 $\pm$ 0.001 0.203 $\pm$ 0.003
AAF (-Attribution) 0.520 $\pm$ 0.009 12.896 $\pm$ 0.002 0.206 $\pm$ 0.004
AAF (+Detect Only) 0.520 $\pm$ 0.009 12.896 $\pm$ 0.002 0.206 $\pm$ 0.004
AAF (+Patch Only) 0.411 $\pm$ 0.007 12.918 $\pm$ 0.001 0.203 $\pm$ 0.003
AAF (+Shaping Only) 0.520 $\pm$ 0.009 12.896 $\pm$ 0.002 0.206 $\pm$ 0.004
50 AAF (Full) 0.626 $\pm$ 0.005 2.475 $\pm$ 0.001 0.156 $\pm$ 0.002
AAF (-Attribution) 0.754 $\pm$ 0.007 2.449 $\pm$ 0.001 0.141 $\pm$ 0.002
AAF (+Detect Only) 0.754 $\pm$ 0.007 2.449 $\pm$ 0.001 0.141 $\pm$ 0.002
AAF (+Patch Only) 0.627 $\pm$ 0.005 2.475 $\pm$ 0.001 0.156 $\pm$ 0.002
AAF (+Shaping Only) 0.754 $\pm$ 0.007 2.449 $\pm$ 0.001 0.141 $\pm$ 0.002
100 AAF (Full) 0.634 $\pm$ 0.005 1.173 $\pm$ 0.001 0.167 $\pm$ 0.002
AAF (-Attribution) 0.706 $\pm$ 0.007 1.159 $\pm$ 0.001 0.159 $\pm$ 0.002
AAF (+Detect Only) 0.706 $\pm$ 0.007 1.159 $\pm$ 0.001 0.159 $\pm$ 0.002
AAF (+Patch Only) 0.630 $\pm$ 0.005 1.174 $\pm$ 0.001 0.169 $\pm$ 0.002
AAF (+Shaping Only) 0.708 $\pm$ 0.008 1.158 $\pm$ 0.002 0.158 $\pm$ 0.003
Compromise–inequality trade-off.
Figure 3 visualizes the Pareto frontier between executed compromise and mean allocation inequality (Gini) for $N=50$ in the canonical setting. AAF interventions shift the system left (fewer compromises) but can increase inequality; penalty_factor and $\alpha$ act as tuning knobs for this trade-off.
Figure 2: Executed compromise ratio by baseline (resource_sharing, $N=50$, canonical setting, $\rho=0$).
Figure 3: Trade-off between executed compromise and mean allocation Gini across baselines (resource_sharing, $N=50$, canonical setting, $\rho=0$).
Figure 4: Mean allocation Gini vs. norm-penalty level for AAF-full (resource_sharing, $N=50$, $\alpha=1.0$, $\rho=0$). Bars report mean $\pm$ 95% CI over 10 seeds.
Byzantine detection and attribution.
Under adversarial injections ($\rho\in\{0.05,0.10\}$, $t_{0}=200$), AAF triggers alarms and assigns responsibility. Across all adversarial runs, the median detection delay is 71 steps (IQR 39–177); Figure 5(a) reports the full CDF on resource_sharing. In the 10% Byzantine setting, AAF achieves mean top-1 attribution accuracy of 0.97 (conditional on a non-missing attribution score), as shown in Figure 5(b).
Figure 5: Detection and Attribution Performance on resource_sharing. (a) Empirical CDF of detection delay under Byzantine injections ($\rho\in\{0.05,0.10\}$, $t_{0}=200$). The median delay is 71 steps. (b) Attribution accuracy (Top-1, Recall@3, Recall@5) across Byzantine fractions $\rho$. Error bars denote 95% CI.
Figure 6: Learning curves under Byzantine injection (resource_sharing, $N=50$, $\rho=0.05$, $t_{0}=200$). Solid lines show the median across 10 seeds with IQR shading for executed compromise rate (top), allocation Gini (middle), and social welfare (bottom). The dashed vertical line marks attack start and the dotted line marks the median first-alarm time (AAF).
7.4 Scaling Analysis
To probe scalability beyond the main factorial grid ($N\leq 100$), we run an additional scaling sweep on resource_sharing with $N\in\{10,50,100,200,500\}$ and $\rho\in\{0,0.05\}$ under the canonical setting ($T=2000$, penalty_factor=0.2, $\alpha=1.0$, full observability). Figure 7(a) shows that AAF’s compromise reduction persists as $N$ grows. At $N=200$, AAF reduces executed compromise by 6.2% ($\rho=0$) and 9.3% ($\rho=0.05$) relative to PPO-only; at $N=500$ the relative reduction is 1.5% and 4.1%, respectively, while welfare remains comparable. Figure 7(b) and Figure 7(c) report wall-clock runtime and communication cost; both scale approximately linearly in $N$, consistent with Proposition 2.
Figure 7: Scaling Analysis on resource_sharing under canonical settings ($\lambda=0.20$, $\alpha=1.0$). (a) Compromise ratio vs. network size (NN). AAF maintains its advantage at scale. (b) Runtime overhead vs. network size. AAF’s computational cost scales linearly. (c) Communication bandwidth vs. network size. AAF’s ledger and attribution messages scale linearly with the number of agents.
7.5 Public Goods Benchmark
To demonstrate the generality of the framework, we evaluate AAF on the public_goods environment, where agents must balance individual contributions against collective returns. Table 3 reports the main results across scales. Consistent with the resource sharing domain, AAF-full achieves the lowest executed compromise ratio among all learning baselines across all scales ($N=10,50,100$), while maintaining or slightly improving social welfare.
Table 3: Results on the Public Goods environment. Values are mean $\pm$ standard error across all hyperparameter regimes and seeds. Bold indicates best among learning baselines (excluding Static Guard). $\downarrow$ = lower is better; $\uparrow$ = higher is better.
NN Baseline Compromise $\downarrow$ Social Welfare $\uparrow$ Gini (Reward) $\downarrow$ Det. Delay
10 AAF (Full) 0.160 $\pm$ 0.000 1.450 $\pm$ 0.001 0.072 $\pm$ 0.000 77 [57–686]
PPO Only 0.196 $\pm$ 0.001 1.440 $\pm$ 0.001 0.073 $\pm$ 0.000
Constrained PPO 0.194 $\pm$ 0.001 1.442 $\pm$ 0.001 0.073 $\pm$ 0.000
Fair PPO 0.196 $\pm$ 0.001 1.440 $\pm$ 0.001 0.073 $\pm$ 0.000
Static Guard 0.000 $\pm$ 0.000 1.577 $\pm$ 0.002 0.073 $\pm$ 0.000
50 AAF (Full) 0.268 $\pm$ 0.001 1.337 $\pm$ 0.001 0.071 $\pm$ 0.000 69 [51–98]
PPO Only 0.311 $\pm$ 0.002 1.325 $\pm$ 0.001 0.073 $\pm$ 0.000
Constrained PPO 0.294 $\pm$ 0.002 1.339 $\pm$ 0.001 0.073 $\pm$ 0.000
Fair PPO 0.311 $\pm$ 0.002 1.325 $\pm$ 0.001 0.073 $\pm$ 0.000
Static Guard 0.000 $\pm$ 0.000 1.562 $\pm$ 0.002 0.070 $\pm$ 0.000
100 AAF (Full) 0.241 $\pm$ 0.001 1.375 $\pm$ 0.001 0.076 $\pm$ 0.000 48 [40–67]
PPO Only 0.271 $\pm$ 0.002 1.363 $\pm$ 0.001 0.077 $\pm$ 0.000
Constrained PPO 0.258 $\pm$ 0.002 1.373 $\pm$ 0.001 0.078 $\pm$ 0.000
Fair PPO 0.271 $\pm$ 0.002 1.363 $\pm$ 0.001 0.077 $\pm$ 0.000
Static Guard 0.000 $\pm$ 0.000 1.550 $\pm$ 0.002 0.068 $\pm$ 0.000
Figure 8 visualizes the compromise reduction across baselines. The relative safety advantage of AAF is particularly pronounced at $N=50$, where it reduces the executed compromise ratio from 0.311 (PPO-only) to 0.268. Notably, the reward Gini coefficient remains stable across all baselines, indicating that the norm enforcement mechanisms do not inadvertently create severe reward disparities in this domain.
Figure 8: Executed compromise ratio by baseline across scales on the public_goods environment ($\rho=0.0$). Bars report mean $\pm$ standard error.
Under adversarial conditions, AAF detects Byzantine injections in the public goods setting with a median delay of 65 steps (IQR 47–93), demonstrating that the sequential hypothesis testing mechanism transfers effectively to different reward structures and interaction topologies without requiring domain-specific tuning.
@@ 7.6 Qualitative Case Study
We also examined representative trajectories to validate that AAF’s numerical improvements correspond to interpretable behavioral shifts. In a typical resource_sharing episode ($N=50$, $T=2000$, $\texttt{penalty\_factor}=0.2$, $\alpha=1.0$, full observability), a coalition of agents repeatedly issued near-maximal requests, producing a visible rise in executed compromises. AAF’s adaptive CUSUM detector raised an alarm shortly after the coalition’s behavior became statistically distinguishable from the baseline, and the intervention pipeline (reward shaping + targeted patching) reduced the coalition’s marginal incentive to over-request. Within a few hundred steps, the episode transitioned back toward a lower-compromise regime. The logged responsibility graph provides a post-hoc audit trail: agents ranked as most responsible typically coincide with injected Byzantine agents when $\rho>0$ (see §7.3).
@@ 7.7 Reproducibility Checklist
  • Code & data: Public GitHub repository: https://github.com/alqithami/AAF. The accompanying artifact bundle includes the aggregated CSVs and the scripts used to regenerate all tables and figures.
  • Environment: Python (tested with 3.12) and pinned dependencies in requirements.txt.
  • Running experiments: scripts/make_grid.py generates JSONL grids; scripts/run_grid.py executes them (supports multi-process runs and sharding via --shard_id / --num_shards); scripts/aggregate.py produces analysis/all_runs_flat.csv and analysis/final_summary.csv; scripts/make_latex.py, scripts/make_figures.py, and scripts/stats.py generate paper-ready assets.
  • Randomness: Each run records its seed in config.json; paired tests are seed-matched.
Empirical evidence supports the analytical intuition: across the full grid, AAF reduces executed compromise relative to PPO in 96% of regimes (median reduction 11.9%), preserves welfare (median change 0.4%), and detects adversarial norm shifts with median delay 71 steps.
@@ 8.1 Why the Results Matter
Because detection and intervention are online, AAF reacts on the same time scale as the norm shift itself. In our Byzantine-injection experiments ($t_{0}=200$), the median detection delay is 71 steps (IQR 39–177). After intervention, the executed-compromise rate typically decreases rather than persisting at the elevated post-attack level, supporting the case for continuous accountability in fast time-scale domains where retrospective audits may arrive too late. This empirical finding strongly supports the theoretical arguments for real-time observability recently advanced by Solomon and others [60].
@@
The bounded-compromise theorem (Thm. 5) predicted that a local, Shapley-weighted penalty would globally stabilize the system once $\lambda H>g_{\max}$. Empirically, we observed $\eta_{\text{emp}}$ within one standard error of the analytic ceiling, even when 5% of agents behaved adversarially. This corroborates the long-standing conjecture that potential-based shaping [43] scales from single- to multi-agent settings when augmented with causal attribution, aligning with recent advances in causal credit assignment [65] and constrained policy optimization for safe RL [1].
@@ 8.2 Limitations and Threats to Validity
Our main factorial grid spanned $\{10,50,100\}$ agents and three penalty magnitudes, and we additionally ran a scaling sweep up to $N=500$ in resource_sharing. A 5000-agent swarm, or penalty factors close to unity, might stress the message bus or saturate the intervention budget. Future work should stress-test the framework under orders-of-magnitude larger populations and adversaries who actively spoof causal edges.
@@ 8.3 Ethical, Legal, and Social Implications
Reward penalties can skew agents toward risk-averse behavior, potentially lowering innovation in financial-market MAS [16]. A human-in-the-loop override—triggered by the yellow-flag dashboard—helps balance safety with exploration.
@@
By design, Shapley-style attribution apportions blame according to marginal causal contribution. If agents differ in observability or computing power, this may correlate with protected attributes, leading to distributive inequity [38]. Auditors should therefore scrutinize $\rho_{i}(e)$ distributions for disparate impact.
@@ 8.4 Relevance to Emerging AI Governance
The EU AI Act (2024) and the NIST AI Risk-Management Framework (2023) both highlight collective emergent risks —but offer few technical recipes for tracing responsibility in networked AI systems. AAF operationalizes these policy aspirations by (i) anchoring an immutable event ledger; (ii) providing real-time statistical tests with formal $\alpha$-level guarantees [27]; and (iii) linking those tests to proportionate, auditable interventions. By addressing the core pillars of responsibility—reliability, transparency, and accountability [15]—and providing a concrete implementation of trust and risk management principles [53], the framework can serve as a reference architecture for regulators drafting sector-specific codes of practice.
@@ 8.5 Future Directions
While the Grafana DAG explorer provides low-level proofs, policy operators may prefer natural-language rationales. Coupling the ledger to an LLM-based explainer guarded by the event digests to prevent hallucination could bridge the accountability-interpretability gap [30].
@@ A.3 Proof of Theorem 3
Under $H_{0}$ the $m$-lag Granger residual–variance ratio has exact $F$ distribution F(m,nm)/mF(m,n{-}m)/m with n=m+1n=m+\!1 degrees-of-freedom [35, 26]. Let $U_{t}\!=\!\mathbf{1}\{F_{t}>h_{t}\}$. Then, by the Chernoff bound for $\chi^{2}$ tails, $\mathbb{P}[U_{t}=1]\leq e^{-h_{t}^{2}/2}=e^{-h_{0}^{2}/2}\,t^{-1}.$ Taking the union bound over $t\geq 1$ yields
@@ A.4 Proof of Theorem 4
Define $f(h)=\mathbb{P}_{H_{0}}(S\geq h)-\alpha$. The process $\{h_{t}\}$ satisfies the Robbins-Monro recursion $h_{t+1}=h_{t}-\eta_{t}f(h_{t})+\eta_{t}\epsilon_{t+1}$ with martingale difference $\epsilon_{t+1}=Z_{t+1}-\mathbb{P}_{H_{0}}(S\geq h_{t})$. Because $S_{t}$ has bounded increments, $\epsilon_{t+1}$ is square integrable with $\mathbb{E}\,\epsilon_{t+1}=0$ and $\sup_{t}\mathbb{E}\,\epsilon_{t+1}^{2}\!<\infty$. Assumption $\sum_{t}\eta_{t}^{2}<\infty$ ensures the Kushner-Clark condition, so $h_{t}\!\to\!h^{\star}$ a.s. where $f(h^{\star})=0$ [31]. Ergodicity of the CUSUM statistic under $H_{0}$ implies $\frac{1}{T}\sum_{t=1}^{T}\mathbf{1}\{S_{t}\geq h^{\star}\}\to\mathbb{P}_{H_{0}}(S\geq h^{\star})=\alpha$, and dominated convergence gives the required limit for $h_{t}$. $\Box$